A normal force 35 N act over an area of 5m2. What is the pressure exerted by the force?


1
Expert's answer
2021-03-08T08:27:00-0500
"P=\dfrac{F}{A}=\dfrac{35\ N}{5\ m^2}=7\ Pa."